OpenStudy (anonymous):

find the exact value of tan40deg. - tan10deg./1+tan40deg.*tan10deg.

OpenStudy (anonymous):

this is in the form of tan(x-y)... \[\tan(40-10)= \frac{1}{\sqrt3}\]

OpenStudy (anonymous):

it is a common trig identity

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

agree

OpenStudy (anonymous):

is this the same as squarert3/3 ??

OpenStudy (anonymous):

\[\frac{\tan (40 {}^{\circ})-\tan (10 {}^{\circ})}{\tan (10 {}^{\circ}) \tan (40 {}^{\circ})+1}=0.57735 \]

OpenStudy (anonymous):

\[\frac{1}{\sqrt3}=\frac{\sqrt{3}}{3}\]

OpenStudy (anonymous):

thank you. I asked that question after you gave me it, I figured that was correct. Thank you!